OpenStudy (anonymous):

Find the slope and y-intercept of each line. f(x)=5/4x+7

OpenStudy (anonymous):

Remember y=mx+b Slope=m Y-intercept=b With that in mind, what do you think the answer is?

OpenStudy (anonymous):

is it ; slope = 5/4 y-intercept = 7

OpenStudy (anonymous):

Yes. You were simply over thinking it. If you remember the formula, this will be easy to you.

OpenStudy (anonymous):

alright , thank you so much! (:

OpenStudy (anonymous):

No problem. :)